find the y-intercept of the line that passes through (1.5,-3.5) and has a slope 3
Answer by lwsshak3(11628)

find the y-intercept of the line that passes through (1.5,-3.5) and has a slope 3
equation of a line: y=mx+b, m=slope, b=y-intercept
For given problem:
Equation: y=3x+b
Using coordinates to find b:
-3.5=3*(1.5)+b
b=-3.5-4.5=-8
y-intercept:-8